OpenStudy (anonymous):

What is the volume , in cubic centimers of a cylinder with a height of 15 cm and a diameter of 12 cm ? a. 180(pi) b. 540(pi) c. 675(pi) d. 2160(pi) I dont know how to answer a question liek that with pi in it . I just know how to solve the equation altogether and i got 1,696.460 :/

OpenStudy (lgbasallote):

the volume of a cylinder = \(\pi r^2 h\) however r is not given..BUT d is given...and if you recall \(d = 2r\) therefore \(r = \frac{d}{2}\) we replace that into the equation \[V = \pi (\frac{d}{2})^2 h\] now just plug in the values

OpenStudy (anonymous):

so it would be 3.14(6)^215) which is 540(pi) which is b ?

OpenStudy (lgbasallote):

yep that's right

OpenStudy (anonymous):

thanks @lgbasallote